card is drawn from an ordinary deck. Find the probability of getting a 3 or a diamond card
1
Expert's answer
2021-10-26T10:18:15-0400

There are 13 diamond cards, four 3's and one diamond 3.

"P(3\; or\; diamond)=\frac{13+4-1}{52}=\frac{16}{52}=\frac{4}{13}\approx0.3077."